Prove that the equation of tangent at the point (at^2 , 2at) on the parabola y^2 = 4ax is x - ty + at^2 = 0
1
Expert's answer
2021-02-01T03:07:31-0500

Here the parabola given in the question is "y^2=4ax" . So, to find the slope at the point "(at^2,2at)" on the parabola we differentiate:-



"y^2=4ax\\"

Differentiating both sides:-



"2y\frac{dy}{dx}=4a\\
\Rightarrow \frac{dy}{dx}=\frac{4a}{2y}=\frac{2a}{y}"

So, slope at the required point:-



"\frac{dy}{dx}|_{(at^2,2at)}=\frac{2a}{2at}=\frac{1}{t}"

So, we know the point and the slope of the curve at that point. So the equation of the tangent is:-



"(y-2at)=\frac{1}{t}(x-at^2)\\~\\
\Rightarrow yt-2at^2=x-at^2\\
\Rightarrow x-ty+at^2=0"

Therefore, the equation of tangent at "(at^2,2at)" on the parabola "y^2=4ax" is:-



"\fcolorbox{red}{cyan}{$\textcolor{black}{x-ty+at^2=0}$}"